# Learning Graph Quantized Tokenizers for Transformers

### v1

@@ Abstract
Transformers serve as the backbone architectures of Foundational Models, where a domain-specific tokenizer helps them adapt to various domains. Graph Transformers (GTs) have recently emerged as a leading model in geometric deep learning, outperforming Graph Neural Networks (GNNs) in various graph learning tasks. However, the development of tokenizers for graphs has lagged behind other modalities, with existing approaches relying on heuristics or GNNs co-trained with Transformers. To address this, we introduce GQT (Graph Quantized Tokenizer), which decouples tokenizer training from Transformer training by leveraging multi-task graph self-supervised learning, yielding robust and generalizable graph tokens. Furthermore, the GQT utilizes Residual Vector Quantization (RVQ) to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities. By combining the GQT with token modulation, a Transformer encoder achieves state-of-the-art performance on 16 out of 18 benchmarks, including large-scale homophilic and heterophilic datasets.
@@ 1 Introduction
Following the success of Transformers (Vaswani et al., 2017) in natural language processing (Devlin et al., 2019; Brown et al., 2020) and computer vision (Dosovitskiy et al., 2021), Graph Transformers (GTs) (Dwivedi & Bresson, 2020; Ying et al., 2021a; Rampášek et al., 2022; Ma et al., 2023; Shirzad et al., 2023; Kong et al., 2023b; Chen et al., 2023; Wu et al., 2022b) have emerged as strong models in geometric deep learning. Unlike message-passing Graph Neural Networks (GNNs), which rely on strong locality inductive biases (Battaglia et al., 2018; Veličković et al., 2018; Hou et al., 2020; Hamilton et al., 2017a; Kipf & Welling, 2017), GTs are inherently more expressive due to their ability to capture long-range interactions between nodes (Ma et al., 2023). This is particularly beneficial in heterophilous settings where local alignment does not hold (Fu et al., 2024). GTs possess an expressive power at least equivalent to the 2-Weisfeiler-Lehman (WL) isomorphism test (Kim et al., 2022), which is sufficient for most real-world tasks (Zopf, 2022). This surpasses the expressive power of message-passing GNNs, which are limited to the 1-WL test (Ying et al., 2021a). Furthermore, a Transformer with sufficient attention heads can match or exceed the expressive power of a second-order invariant graph network, outperforming message-passing GNNs (Kim et al., 2022). However, both GNNs and Transformers are susceptible to over-smoothing (Li et al., 2018; Zhou et al., 2021; Dovonon et al., 2024).
@@
GTs require consideration of both graph structure and features, as nodes with identical features will otherwise be projected into the same representation regardless of their surrounding structures (Hoang et al., 2024). There are three general approaches to address this limitation (Hoang et al., 2024): (1) node feature modulation, which involves injecting structural information into the node features; (2) context node sampling, where a sampling strategy is used to construct a sequence over the neighbor nodes; and (3) modifying the architecture of a vanilla Transformer to directly incorporate structural biases. Given that Transformers are universal approximators of sequence-to-sequence functions (Yun et al., 2020) and considering the rapid developments in efficient implementation of multi-head attention (MHA) module (Dao et al., 2022a; Liu et al., 2024), which enables longer context sizes of up to million-scale tokens (Reid et al., 2024), a well-designed graph tokenizer can allow a vanilla Transformer model to efficiently process even large-scale graphs. Recent studies on applying Large Language Models (LLMs) to graph-related tasks have found that representing graphs through textual descriptions can lead to surprisingly strong performance gains that surpass those of GNNs, suggesting that vanilla Transformers are indeed capable of effectively learning graph structures (Ye et al., 2024; He et al., 2024). Nonetheless, LLMs are not inference-efficient, and hence our goal in this paper is to devise a lightweight and efficient graph tokenization strategy that enables vanilla Transformer encoders to learn graph structures effectively, without relying on LLMs.
@@
Tokenizers typically employ self-supervised objectives to abstract data into a sequence of discrete tokens, enabling Transformers to learn representations across various modalities as a unified stream of data. This discretization is achieved through vector quantization techniques (Van Den Oord et al., 2017; Lee et al., 2022), which offer several benefits, including: (1) significantly reduced memory requirements, (2) improved inference efficiency, (3) allowing Transformers to focus on long-range dependencies rather than local information, and (4) the capacity to learn more high-level representations due to a compact latent space (Yuan et al., 2021; Yu et al., 2022). These advantages are particularly important in auto-regressive generative modeling, where quantized tokens allow Transformers to generate high-quality outputs in multiple modalities (Dubey et al., 2024; Lee et al., 2022; Dhariwal et al., 2020; Ramesh et al., 2021; Team, 2024). Despite its importance in other domains, tokenization remains under-explored for graph-structured data. To address this limitation, we propose the Graph Quantized Tokenizer (GQT), a novel approach that learns a hierarchical sequence of tokens over graphs using self-supervised objectives tailored to graph-structured data. More specifically, our contributions are as follows:
@@
  • We propose a graph tokenizer that utilizes multi-task graph self-supervised objectives to train a graph encoder, enabling it to fully capture local interactions and allowing the Transformer to focus on long-range dependencies.
@@
  • We introduce a novel combination of semantic edges and random walks to facilitate the Transformer’s access to long-range interactions, and employ hierarchical encoding and gating mechanisms to modulate the tokens and provide informative representations to the Transformer.
  • Through extensive experiments on both homophilic and heterophilic datasets, including large-scale benchmarks, we demonstrate that our proposed tokenizer enables a Transformer encoder to achieve state-of-the-art performance on 16 out of 18 benchmarks while substantially reducing the memory footprint of the embeddings.
@@ 2 Related Works
Graph Transformers (GTs) have shown promising performance on various graph learning tasks, surpassing GNNs on many benchmarks. GTs can be broadly categorized into two directions (Hoang et al., 2024; Müller et al., 2024): (1) modifying the vanilla Transformer architecture to incorporate structural inductive biases, or (2) encoding the input graph to make it compatible with the vanilla Transformer design. Early examples of the first approach include Graph Attention Network (Veličković et al., 2018), which uses an attention module to compute pairwise node attentions and masks the attention matrix based on connectivity information. Subsequent works have replaced the scaled-dot attention module with various structure-aware sparse attention modules (Rampášek et al., 2022; Bo et al., 2023; Ying et al., 2021a; Deng et al., 2024; Wu et al., 2023b; Liu et al., 2023a; Chen et al., 2022; Dwivedi & Bresson, 2020; Shirzad et al., 2023; Ma et al., 2023). Examples of the second approach include Graph Memory Network (Khasahmadi et al., 2020), which passes non-linear projections of node features and structural encoding to a Transformer-like model. Structural encoding methods, such as Laplacian eigenvectors or Random walk-based encoding (Dwivedi et al., 2022; Ma et al., 2023; Cantürk et al., 2024), allow injecting structural information directly into the node features. Another approach involves using GNNs to encode local structure along with node features, followed by passing the representation to vanilla Transformers to capture long-range dependencies. (Rong et al., 2020; Wu et al., 2021; Chen et al., 2023; 2022). Recent studies leverage LLMs, where graphs are represented through natural language expressions, and an LLM performs graph-related tasks through in-context learning, instruction tuning, or soft prompts (Fatemi et al., 2024; Ye et al., 2024; He et al., 2024). For a detailed survey on GTs, see (Müller et al., 2024; Hoang et al., 2024).
@@
Graph Tokenization provides GTs with rich node tokens that encapsulate both structural and semantic information. Various approaches have been proposed to define these node tokens. TokenGT (Kim et al., 2022) treats nodes and edges as independent tokens defined by their features, type identifiers, and structural encodings. NAGphormer (Chen et al., 2023) represents each node as a set of $L$ tokens, where the $l^{th}$ token is the representation of the node from $l^{th}$ hop aggregation. In contrast, GraphiT (Mialon et al., 2021) defines a node token as the concatenation of its feature and representation from a graph convolutional kernel network (GCKN). VCR-Graphormer (Fu et al., 2024) expands the notion of node tokens to include sequences comprising the node feature and features of semantically and community-related neighboring nodes. SGT (Liu et al., 2023b) is a non-parametric tokenizer designed for molecular tasks, which simplifies the tokenization process to a non-parametric graph operator without non-linearity, unlike motif-based tokenizers (Zhang et al., 2021; Jin et al., 2018) or GNN pre-training methods (Xia et al., 2023). NodePiece (Galkin et al., 2022) is a knowledge-graph tokenizer that represents a target node as a hash of its top-k closest anchors, their distances, and relational context. For a more detailed review see Müller et al. (2024). While vector quantization (VQ) (Van Den Oord et al., 2017; Lee et al., 2022) has been explored in other modalities, its application in graph learning is limited. Notable exceptions include VQGraph (Yang et al., 2024), which employs VQ for graph distillation, and NID (Luo et al., 2024a), which uses VQ to learn discrete node IDs for downstream prediction tasks.
@@ 3 Preliminaries
Messag-Passing GNNs. Let $\mathcal{G}$ denote the space of graphs. A graph $g$ within this space is defined as $g=\left(\mathcal{V},\mathcal{E},\textbf{X},\textbf{E}\right)$ where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. $\textbf{X}\in\mathbb{R}^{\left|\mathcal{V}\right|\times d_{x}}$ represents the node features of dimension $d_{x}$, and $\textbf{E}\in\mathbb{R}^{\left|\mathcal{V}\right|\times\left|\mathcal{V}\right|\times d_{e}}$ represents the edge features of dimension $d_{e}$. A message-passing GNN takes $g$ as input and learns representations $h^{l}_{v}$ for $v\in\mathcal{V}$ ($h^{0}_{v}=x_{v}$) in each layer $l$ as follows (Gilmer et al., 2017):
$h^{l}_{v}=f^{l}_{\theta}\left(h^{l-1}_{v},g^{l}_{\phi}\left(\left\{\left(h^{l-1}_{v},h^{l-1}_{u},e_{uv}\right)|u\in\mathcal{N}_{i}(v)\right\}\right)\right)$ (1)
where $f_{\theta}$ and $g_{\phi}$ are known as update (combine) and message (aggregate) functions, respectively, and $\mathcal{N}_{i}(v)$ denotes the set of immediate neighbors of the node $v$. With this representation, we can perform various tasks, including node classification as $\text{MLP}\left(h_{v}\right)$, edge prediction as $\text{MLP}\left(h_{u}\odot h_{v}\right)$, or graph classification as $\text{MLP}\left(\mathcal{R}\left(\left\{h_{u}|u\in\mathcal{V}\right\}\right)\right)$ where $\mathcal{R}$ is a pooling (readout) function.
Graph Transformers, on the other hand, first use a tokenizer $T_{v}=\mathcal{T}_{\psi}\left(\mathcal{N}(v)\right)$ to map each node $v\in\mathcal{V}$ into a sequence of tokens $T_{v}$ by considering some notion of neighborhood $\mathcal{N}$. The simplest design is when $\mathcal{N}$ is zero-hop neighborhood (i.e., the node itself) and $\mathcal{T}_{\psi}$ is a node feature lookup function. The neighborhood $\mathcal{N}$ can be extended to include an ego network (Zhao et al., 2021) or top-k Random Walk based neighbors (Fu et al., 2024), and $\mathcal{T}_{\psi}$ can be enhanced to representations from a GNN (Chen et al., 2023). Once the tokens are computed, along with a node positional encoding function (PE), we can define the input to a Transformer as $h^{0}_{v}=\left[T_{v}||\text{PE}\left(v\right)\right]$ and compute the representation in each layer $l$ of a vanilla Transformer encoder as follows:
@@
Vector Quantization projects embeddings $\textbf{X}\in\mathbb{R}^{n\times d_{x}}$ into a more compact space of codebooks $\textbf{C}\in\mathbb{R}^{k\times d_{c}}$, where $k\ll n$. The codebooks can be learned by minimizing various objectives such as K-means clustering. The new representation of $x_{i}$ is then computed as follows (Van Den Oord et al., 2017):
$z(x_{i})=c_{k}\quad\text{where}\quad k=\operatorname*{arg\,min}_{j}\|x_{i}-c_{j}\|_{2}^{2}$ (4)
Building upon this concept, RQ-VAE (Lee et al., 2022) extends VQ to a sequence of codebooks, where each consecutive codebook quantizes the residual error from the previous codebook, i.e., $r_{i}=z_{i}-c_{k}$. This hierarchical approach constructs a multi-level quantized representation, enhancing the overall quantization quality.
@@ 4.1 Tokenizer Properties
Our goal is to design a graph tokenizer that can learn to generate tokens that exhibit three key characteristics, which are essential for effective graph representation learning. These characteristics are as follows.
@@
Local Interactions. The learned tokens should encapsulate local interactions, allowing the Transformer to focus on global dependencies. This is analogous to Vision Transformers (ViTs), where the Transformer attends to image patches instead of pixels, enabling efficient learning on abstract tokens (Dosovitskiy et al., 2021; Liu et al., 2021). To achieve a similar effect on graph-structured data, we leverage message-passing GNNs as the foundation of the tokenizer’s encoder, capitalizing on their strong locality inductive bias to effectively capture local interactions in the representation space (Battaglia et al., 2018). Our design accommodates various GNN layer choices without constraints; for simplicity, we opt for the widely used Graph Attention Network (GAT) (Veličković et al., 2018) as our base graph encoder. The representation of node $i$ in layer $l$ is computed as:
$h_{i}^{l}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\textbf{W}h_{j}^{(l-1)}\right),\quad\alpha_{ij}=\frac{\exp\left(\sigma\left(\textbf{W}_{2}\left[\textbf{W}_{1}h_{i}^{(l-1)}\|\textbf{W}_{1}h_{j}^{(l-1)}\right]\right)\right)}{\sum\limits_{k\in\mathcal{N}(i)}\exp\left(\sigma\left(\textbf{W}_{2}\left[\textbf{W}_{1}h_{i}^{(l-1)}\|\textbf{W}_{1}h_{k}^{(l-1)}\right]\right)\right)}$ (5)
where $\sigma$ is a non-linearity, and $\alpha_{ij}$ is the normalized attention score between two connected nodes $i$ and $j$.
@@
Memory Efficiency. The tokens should be compact to facilitate efficient memory usage, enabling the Transformer to perform efficient inference. To achieve this, we introduce a Residual-VQ (RVQ) (Lee et al., 2022) layer to quantize the GNN representations into a sequence of discrete tokens. Quantization not only helps with generalization due to its regularization effect but also significantly reduces memory usage. Using an RVQ with $c$ codebooks (typically $c=\{2,\cdots,8\}$), a graph with feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{x}}$ can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{N\times c}$ and codebook representation of $\textbf{C}\in\mathbb{R}^{c\times K\times d_{c}}$, where $c$ is the number of codebooks (i.e., levels of quantization), $K$ is the codebook size, and $d_{c}$ is the code dimension. To illustrate the benefits of this approach, consider a graph with $10^{6}$ nodes and a feature dimension of 1024 ($\textbf{X}\in\mathbb{R}^{10^{6}\times 1024}$). Using an RVQ with 3 codebooks and a codebook size of 256, this graph can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{10^{6}\times 3}$ plus $\textbf{C}\in\mathbb{R}^{3\times 256\times 1024}$, resulting in a 270-fold reduction in required memory.
@@
Robustness and Generalization. The tokens should be robust and generalizable. To achieve this, we rely on graph self-supervised learning. Self-supervised representations have been shown to be more robust to class imbalance (Liu et al., 2022) and distribution shift (Shi et al., 2023), while also capturing better semantic information (Assran et al., 2023) compared to representations learned through supervised objectives. Moreover, self-supervised graph representations have demonstrated superior performance on downstream tasks compared to representations learned in a fully supervised manner, indicating better generalization capabilities (Hu et al., 2020b; Sun et al., 2020; You et al., 2020; 2021; Hassani & Khasahmadi, 2020; Hou et al., 2022; Veličković et al., 2019; Zhu et al., 2020b; Thakoor et al., 2022). Additionally, multi-task learning with self-supervised objectives has been shown to achieve better performance on downstream tasks (Doersch & Zisserman, 2017; Ghiasi et al., 2021). To leverage these benefits, we propose training the GNN encoder with three self-supervised objectives. Unlike RQ-VAE (Lee et al., 2022), which uses reconstruction as its primary objective, we employ graph-specific objectives to capture the nuances of both structure and features within the tokens. Specifically, we use Deep Graph Infomax (DGI) (Veličković et al., 2019) and Graph Masked Auto-Encoder 2 (GMAE2) (Hou et al., 2023). DGI is a contrastive method that contrasts local (node) encoding with global (graph or sub-graph) encoding, whereas GMAE2 combines generative and distillation objectives to jointly reconstruct masked features and track teacher representations.

### v2

@@ Abstract
Transformers serve as the backbone architectures of Foundational Models, where domain-specific tokenizers allow them to adapt to various domains. Graph Transformers (GTs) have recently emerged as leading models in geometric deep learning, outperforming Graph Neural Networks (GNNs) in various graph learning tasks. However, the development of tokenizers for graphs has lagged behind other modalities. To address this, we introduce GQT (Graph Quantized Tokenizer), which decouples tokenizer training from Transformer training by leveraging multi-task graph self-supervised learning, yielding robust and generalizable graph tokens. Furthermore, the GQT utilizes Residual Vector Quantization (RVQ) to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities. By combining the GQT with token modulation, a Transformer encoder achieves state-of-the-art performance on 20 out of 22 benchmarks, including large-scale homophilic and heterophilic datasets.
@@ 1 Introduction
Following the success of Transformers (Vaswani et al., 2017) in Natural Language Processing (Devlin et al., 2019; Brown et al., 2020) and Computer Vision (Dosovitskiy et al., 2021), Graph Transformers (GTs) (Dwivedi and Bresson, 2020; Ying et al., 2021; Rampášek et al., 2022; Shirzad et al., 2023; Chen et al., 2023; Wu et al., 2022) have emerged as strong models in geometric deep learning. In contrast to message-passing Graph Neural Networks (GNNs), which are inherently constrained by strong locality inductive biases (Battaglia et al., 2018; Veličković et al., 2018; Hou et al., 2020; Hamilton et al., 2017; Kipf and Welling, 2017; Zheng et al., 2024), GTs exhibit greater expressivity due to their capacity to capture long-range interactions between nodes (Ma et al., 2023; Kim et al., 2022; Zopf, 2022). This is particularly beneficial in heterophilic settings where local alignment does not hold (Fu et al., 2024). This dichotomy highlights a fundamental trade-off between GNNs, which focus on local neighborhood aggregation, and GTs, which employ pairwise attention to model global graph structures. A natural question arises: can we synergistically integrate the strengths of both approaches to leverage the complementary benefits of local and global representations? Specifically, is it possible to harness the locality-aware representations learned by GNNs to construct discrete tokens, thereby enabling GTs to operate efficiently while still capturing salient graph properties?
@@
GTs require consideration of both graph structure and features, as nodes with identical features will otherwise be projected into the same representation regardless of their surrounding structures (Hoang et al., 2024). There are three general approaches to address this limitation (Hoang et al., 2024): (1) node feature modulation, which involves injecting structural information into the node features; (2) context node sampling, where a sampling strategy is used to construct a sequence over the neighbor nodes; and (3) modifying the architecture of a vanilla Transformer to directly incorporate structural biases. Given that Transformers are universal approximators of sequence-to-sequence functions (Yun et al., 2020) and considering the rapid developments in efficient implementation of Multi-Head Attention (MHA) module (Dao et al., 2022; Liu et al., 2024), which enables longer context sizes (Reid et al., 2024), we argue that a well-designed graph tokenizer can allow a vanilla Transformer to efficiently process even large-scale graphs. Recent studies on applying Large Language Models (LLMs) to Text-Attributed Graphs (TAGs) have shown surprisingly strong performance gains surpassing those of GNNs, suggesting that vanilla Transformers are indeed capable of effectively learning graph structures (Ye et al., 2024; Xu et al., 2025). Nonetheless, LLMs are not efficient at inference time. Our goal is to devise a lightweight and efficient graph tokenizater that allows vanilla Transformer encoders to learn graph structures effectively.
@@
Tokenizers typically employ self-supervised objectives to abstract data into a sequence of discrete tokens, allowing Transformers to learn representations across various modalities as a unified stream of data. The discretization is usually achieved through vector quantization techniques (Van Den Oord et al., 2017; Lee et al., 2022), which offer several benefits, including: (1) significantly reduced memory requirements, (2) improved inference efficiency, (3) allowing Transformers to focus on long-range dependencies rather than local information, and (4) the capacity to learn more high-level representations due to a compact latent space (Yuan et al., 2021; Yu et al., 2022a). These advantages are particularly important in auto-regressive generative modeling, where quantized tokens allow Transformers to generate high-quality outputs in multiple modalities (Dubey et al., 2024; Lee et al., 2022; Dhariwal et al., 2020; Ramesh et al., 2021; Team, 2024). Despite its importance in other domains, tokenization remains under-explored for graph-structured data. To address this limitation, we propose the Graph Quantized Tokenizer (GQT), a novel approach that learns a hierarchical sequence of tokens over graphs using self-supervised objectives tailored to graph-structured data. More specifically, our contributions are as follows:
@@
  • We propose a graph tokenizer that uses multi-task graph self-supervised learning to train a graph encoder, enabling it to fully capture local interactions and allowing the Transformer to focus on long-range dependencies.
@@
  • We introduce a novel combination of semantic edges and random walks to facilitate access to long-range interactions, and employ hierarchical encoding and gating mechanisms to modulate the tokens and provide informative representations to the Transformer.
  • Through extensive experiments on both homophilic and heterophilic datasets, including large-scale and long-range benchmarks, we demonstrate that our tokenizer enables Transformer encoders to achieve state-of-the-art performance on 20 out of 22 benchmarks while substantially reducing the memory footprint of the embeddings.
@@ 2 Related Works
Graph Transformers (GTs) have shown promising performance on various graph learning tasks, surpassing GNNs on many benchmarks. Designing GTs can be broadly categorized into two directions (Hoang et al., 2024; Müller et al., 2024): (1) modifying the vanilla Transformer architecture to incorporate structural inductive biases, or (2) encoding the input graph to make it compatible with the vanilla Transformer. Early examples of the first approach include Graph Attention Network (GAT) (Veličković et al., 2018), which uses an attention module to compute pairwise node attention and masks the attention matrix based on connectivity information. Subsequent works have replaced the scaled-dot attention module with various structure-aware sparse attention modules (Rampášek et al., 2022; Bo et al., 2023; Ying et al., 2021; Deng et al., 2024; Wu et al., 2023b; Liu et al., 2023a; Chen et al., 2022; Dwivedi and Bresson, 2020; Shirzad et al., 2023; Ma et al., 2023). Graph Memory Network (GMN) (Khasahmadi et al., 2020) is an example of the second approach, which passes non-linear projections of node features and structural encoding to a Transformer-like model. Structural encodings such as Laplacian eigenvectors or Random walk-based encoding (Dwivedi et al., 2022a; Ma et al., 2023; Cantürk et al., 2024), allow injecting structural information directly into the node features. Some works use GNNs to encode local structure along with node features into embeddings that are passed to vanilla Transformers to capture long-range dependencies (Rong et al., 2020; Wu et al., 2021; Chen et al., 2023, 2022). Recent studies leverage LLMs, where graphs are represented through natural language, and an LLM performs graph-related tasks through in-context learning, instruction-tuning, or soft-prompting (Fatemi et al., 2024; Ye et al., 2024; He et al., 2024). For a detailed survey on GTs, see (Müller et al., 2024; Hoang et al., 2024).
@@
Graph Tokenization provides GTs with rich node tokens that encapsulate both structural and semantic information. TokenGT (Kim et al., 2022) treats nodes and edges as independent tokens defined by their features, type identifiers, and structural encodings. NAGphormer (Chen et al., 2023) represents each node with $L$ tokens, where the $l^{th}$ token is the representation of the node from the $l^{th}$ hop aggregation. GraphiT (Mialon et al., 2021) defines a node token as the concatenation of its feature and representation from a graph convolutional kernel network (GCKN). VCR-Graphormer (Fu et al., 2024) expands the notion of node tokens to include sequences comprising the node feature and features of semantically and community-related neighboring nodes. SGT (Liu et al., 2023c) is a non-parametric tokenizer designed for molecular tasks, which unlike motif-based tokenizers (Zhang et al., 2021; Jin et al., 2018) or GNN pre-training methods (Xia et al., 2023), simplifies the tokenization process to a non-parametric graph operator without non-linearity. NodePiece (Galkin et al., 2022) is a knowledge-graph tokenizer that represents a target node as a hash of its top-k closest anchors, their distances, and relational context. While Vector Quantization (VQ) has been explored in other modalities (Van Den Oord et al., 2017; Lee et al., 2022; Yu et al., 2022b; Van Kempen et al., 2024; Li et al., 2024), its application in graph learning is limited. Notable exceptions include VQ-GNN (Ding et al., 2021), which uses quantized representations combined with a low-rank graph convolution matrix to avoid neighbor explosion problem, VQGraph (Yang et al., 2024), which employs VQ for distilling a GNN into an MLP, and NID (Luo et al., 2024a), which uses VQ to learn discrete node IDs for downstream prediction tasks.
@@ 3 Preliminaries
Messag-Passing GNNs. Let $\mathcal{G}$ denote the space of graphs. A graph $g\in\mathcal{G}$ is defined as $\left(\mathcal{V},\mathcal{E},\textbf{X},\textbf{E}\right)$ where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. $\textbf{X}\in\mathbb{R}^{\left|\mathcal{V}\right|\times d_{x}}$ represents the node features of dimension $d_{x}$, and $\textbf{E}\in\mathbb{R}^{\left|\mathcal{V}\right|\times\left|\mathcal{V}\right|\times d_{e}}$ represents the edge features of dimension $d_{e}$. A message-passing GNN takes $g$ as input and learns representations $h^{l}_{v}$ for $v\in\mathcal{V}$ ($h^{0}_{v}=x_{v}$) in each layer $l$ as follows (Gilmer et al., 2017):
$h^{l}_{v}=f^{l}_{\theta}\left(h^{l-1}_{v},g^{l}_{\phi}\left(\left\{\left(h^{l-1}_{v},h^{l-1}_{u},e_{uv}\right)|u\in\mathcal{N}_{v}\right\}\right)\right)$ (1)
where $f_{\theta}$ and $g_{\phi}$ are known as combine and aggregate functions, respectively. $\mathcal{N}_{v}$ denotes the set of immediate neighbors of the node $v$. Once the node representations are computed, we can perform various tasks including node classification as $\text{MLP}\left(h_{v}\right)$, edge prediction as $\text{MLP}\left(h_{u}\odot h_{v}\right)$, or graph classification as $\text{MLP}\left(\mathcal{R}\left(\left\{h_{u}|u\in\mathcal{V}\right\}\right)\right)$, where $\mathcal{R}$ is a pooling (readout) function.
Graph Transformers use a tokenizer $T_{v}=\mathcal{T}_{\psi}\left(\mathcal{N}(v)\right)$ to map each node $v\in\mathcal{V}$ into a sequence of tokens $T_{v}$ by considering a notion of neighborhood $\mathcal{N}$. The simplest design is when $\mathcal{N}$ is zero-hop neighborhood (i.e., the node itself) and $\mathcal{T}_{\psi}$ is a node feature lookup function. The neighborhood $\mathcal{N}$ can be extended to include the node’s ego network (Zhao et al., 2021) or top-k Random Walk based neighbors (Fu et al., 2024), and $\mathcal{T}_{\psi}$ can be enhanced to representations from a GNN (Chen et al., 2023). Node tokens along with positional encodings (PE) are passed to the Transformer as $h^{0}_{v}=\left[T_{v}||\text{PE}\left(v\right)\right]$. The representations in the $l^{th}$ layer of a Transformer encoder are computed as:
@@
Vector Quantization projects embeddings $\textbf{X}\in\mathbb{R}^{n\times d_{x}}$ into a more compact space of codebooks $\textbf{C}\in\mathbb{R}^{k\times d_{c}}$, where $k\ll n$. The codebooks can be learned by minimizing various objectives such as K-means clustering. The new representation of $x_{i}$ is then computed as (Van Den Oord et al., 2017):
[…]
Building upon this concept, Residual-VQ (RVQ) (Lee et al., 2022) extends VQ to a sequence of codebooks, where each consecutive codebook quantizes the residual error from the previous codebook, i.e., $r_{i}=z_{i}-c_{k}$. This hierarchical approach constructs a multi-level quantized representation, enhancing the overall quantization quality. More details of RVQ are included in Appendix 9.
@@ 4.1 Tokenizer Properties
Our goal is to design a graph tokenizer that learns node tokens that exhibit three key characteristics:
@@
Modeling Local Interactions. The tokens should encapsulate local interactions, allowing the Transformer to focus on long-range dependencies. This is analogous to Vision Transformers (ViTs), where the Transformer attends to image patches instead of pixels (Dosovitskiy et al., 2021; Liu et al., 2021). To achieve this, we leverage GNNs as the tokenizer encoder to model local interactions in the representation space (Battaglia et al., 2018). Our design accommodates various GNN layer choices without constraints; for simplicity, we opt for a GAT encoder (Veličković et al., 2018).
@@
Memory Efficiency. The tokens also should be compact to facilitate efficient memory usage. To achieve this, we introduce a Residual-VQ (RVQ) (Lee et al., 2022) layer to quantize the GNN representations into a sequence of discrete tokens. Quantization not only helps with generalization due to its regularization effect but also significantly reduces memory usage. Using an RVQ with $c$ codebooks (typically $c=\{2,\cdots,8\}$), a graph with feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{x}}$ can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{N\times c}$ and codebook representation of $\textbf{C}\in\mathbb{R}^{c\times K\times d_{c}}$, where $c$ is the number of codebooks (i.e., levels of quantization), $K$ is the codebook size, and $d_{c}$ is the code dimension. To illustrate the benefits of this approach, consider a graph with $10^{6}$ nodes and a feature dimension of 1024 ($\textbf{X}\in\mathbb{R}^{10^{6}\times 1024}$). Using an RVQ with 3 codebooks and a codebook size of 256, this graph can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{10^{6}\times 3}$ plus $\textbf{C}\in\mathbb{R}^{3\times 256\times 1024}$, resulting in a 270-fold reduction in memory.
@@
Robustness and Generalization. The tokens should be robust and generalizable. To achieve this, we rely on graph self-supervised learning. Self-supervised representations have been shown to be more robust to class imbalance (Liu et al., 2022) and distribution shift (Shi et al., 2023), while also capturing better semantic information (Assran et al., 2023) compared to representations learned through supervised objectives. Moreover, self-supervised graph representations have demonstrated superior performance on downstream tasks compared to representations learned in a fully supervised manner, indicating better generalization capabilities (Hu et al., 2020b; Sun et al., 2020; You et al., 2020; Fu et al., 2020; You et al., 2021; Hassani and Khasahmadi, 2020; Veličković et al., 2019; Zhu et al., 2020b). Additionally, multi-task learning with self-supervised objectives has been shown to achieve better performance on downstream tasks (Doersch and Zisserman, 2017; Ghiasi et al., 2021). To leverage these benefits, we propose training the GNN encoder with three self-supervised objectives. Unlike RQ-VAE (Lee et al., 2022), which uses reconstruction as its primary objective, we employ graph-specific objectives to capture the nuances of both structure and features.